OpenStudy (anonymous):

a recipe calls for 2/5 cups of chicken broth and 1/6 cup of vetable broth.what is the total amount of broth need for the recipe

OpenStudy (anonymous):

2/5 * 6/6 = 12/30 1/6 * 5/5= 5/30 12/30 + 5/30 = 17/30

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem (: anything else?

OpenStudy (anonymous):

just a second

OpenStudy (anonymous):

you have 4/6 cup of flour in your cupboard.A recipe for bread calls for 1/4 cup of flour.How much flour would you have left if you made the bread